# Fairness in Ranking under Disparate Uncertainty

### v3

@@ 6. Extension to $G$ Groups
In this section, we discuss the extension of the EOR algorithm beyond two groups. In particular, we consider the general case where a candidate belongs to one of G groups $g\in[1\cdots G]$. From Section 4, we can generalize the cost burden to the principal similar to (4), taking all the groups into account for the normalization factor as follows
@@
(8) $\delta(\sigma)=\max_{g}\left\{\frac{nRel(g|\sigma)}{nRel(g)}\right\}-\min_{g}\left\{\frac{nRel(g|\sigma)}{nRel(g)}\right\}.$
@@
Note that the above selection rule is a strict generalization of Algorithm 1 and it reflects the intuition of minimizing the gap in relevance proportions for all the groups. It can be verified that the runtime complexity with selection rule according to equations (8), (9) for a constant number of groups $G$ is ${O}(n\log{n}+Gn)$. Furthermore, we can extend the cost-approximation guarantee to the multi-group case.
@@
The EOR rankings are cost optimal up to a gap of $\phi\delta_{max}$ for $G$ groups, such that
$\phi$ $=$ $\frac{2\delta(\sigma^{EOR}_{k})}{(G-1)\sum_{g=1}^{G}nRel(g)}\left(\sum_{\{A,B\}}\left|\frac{p_{A}-p_{B}}{q_{A}+q_{B}}\right|\right),$
@@ 7. Experimental Evaluation
We now evaluate the EOR framework and algorithm empirically and compare against several baselines – namely Demographic (Statistical) Parity ($\pi^{DP}$) (Yang and Stoyanovich, 2016), FAIR Ranking Principle ($\pi^{FS}$) (Zehlike et al., 2017), Probability Ranking Principle ($\pi^{PRP}$) (Robertson, 1977), Thompson Sampling Policy ($\pi^{TS}$) (Singh et al., 2021), Uniform Policy ($\pi^{\text{unif}}$), Disparate Treatment of Exposure ($\pi^{EXP}$) (Singh and Joachims, 2018), and Fair Rank Aggregation ($\pi^{RA}$) (Cachel and Rundensteiner, 2023) with proportional representation of exposure. We discuss implementation details of these baselines in Section E.1.
@@ 7.1.1. How does $\pi^{EOR}$ compare against the baselines under varying amounts of disparate uncertainty?
As predicted by theory, $\pi^{EOR}$ maintains low unfairness at all levels of disparate uncertainty, outperforming all the baselines $\pi^{PRP}$, $\pi^{DP}$, $\pi^{TS}$, $\pi^{EXP}$, $\pi^{RA}$, and $\pi^{FS}$. Note that $\pi^{EOR}$ even outperforms the uniform policy $\pi^{\text{unif}}$, since any individual ranking drawn from $\pi^{\text{unif}}$ is likely to be unfair. In terms of effectiveness, the theoretically optimal skyline is given by $\pi^{PRP}$. Across all levels of disparate uncertainty, $\pi^{EOR}$ is at least competitive with the other baselines, indicating that the EOR criterion does not impose a disproportionate cost of fairness for the Principal.
@@ 7.1.2. At which positions in the rankings do the policies incur unfairness?
While the previous table summarized unfairness across the whole ranking, Figure 4 (left) provides more detailed insights into how unfairness accumulates across positions in the ranking. The only method that is systematically fair across all positions $k$ is $\pi^{EOR}$, keeping the unfairness $\delta(\sigma_{k})$ from 4.1 close to zero everywhere in the ranking. The baselines generally start accumulating unfairness towards one group right from the top of the ranking. Their unfairness only decreases once they run out of viable candidates from the group they prefer. The only exception is $\pi^{\text{unif}}$, here for a specific ranking with median $\sum_{k=1}^{n}|\delta(\sigma^{\text{unif}}_{k})|$. However, rankings from $\pi^{\text{unif}}$ tend to stray much further from zero than the $\pi^{EOR}$ ranking. Additional results for the medium and low disparate uncertainty settings in Figure 11 of Section E.2 further support these findings.
Figure 4. Left: EOR fairness $\delta(\sigma_{k})$, Middle: group costs according to (3), Right: the principal’s total cost according to (4) of the ranking policies for the synthetic dataset with high disparate uncertainty shown in top right of Table 1. Group A consists of 30 candidates with sharp probabilities with $p_{i}\sim\text{Beta}(1/20,1/20)$. This provides $nRel(A)=14.96$ expected number of relevant candidates. Group B also has similar candidates, in particular, it has 31 candidates, with relatively flat probabilities $p_{i}\sim\text{Beta}(5,5)$, providing $nRel(B)=14.94$ expected number of relevant candidates.
@@ 7.1.3. How do the ranking policies distribute the costs between the stakeholders?
In Figure 4 (middle) we investigate how the ranking policies distribute the cost $c(g|\pi_{k})$ from Equation (3) between group A and group B. It shows that only $\pi^{EOR}$ has an equal cost to both groups across the whole ranking, which can be seen from the overlapping cost curves for both groups. Furthermore, the cost is substantially lower for both groups than their expected cost under the uniform policy (diagonal line).Figure 4 (right) shows the total cost to the principal, and again $\pi^{EOR}$ is competitive with the baselines.
All other baselines incur substantially disparate costs to the groups, some even worse than the uniform lottery. In particular, $\pi^{DP}$ selects the candidates alternately between the two groups since group sizes are relatively similar, but this results in selecting a higher proportion of relevance from group A because the relevance probabilities are sharper for group A than for B. As a result, the cost burden is higher for group B. $\pi^{TS}$ is fairer than $\pi^{PRP}$, since it randomizes relevant candidates before sorting them in decreasing order of relevance, however being group oblivious, it still places an uneven cost burden.
@@
Figure 5. US Census Dataset: EOR fairness $\delta(\sigma_{k})$ and cost of the ranking policies computed with true relevance labels from the test subset for the US Census dataset. Top: Two groups setting using the White and Black/African American racial groups for the state of Alabama. Bottom: Multiple (four groups) setting using White, Black/African American, Asian, and Other for the state of NY.
@@ 7.2. US Census Survey Data
While the synthetic experiments provide insights into the behavior of the methods under varying conditions, we now investigate how far $\pi^{EOR}$ can mitigate unfairness as it arises in real-world datasets where the relevance probabilities $\mathbb{P}(r_{i}|\mathcal{D})$ are learned from data. In particular, we consider the US Census Survey dataset (Ding et al., 2021) for the year 2018 and the state of Alabama and New York, consisting of 22,268 and 103,021 records respectively. The task is to predict whether the income for an individual >\$$>\$50K$ based on features such as educational attainment, occupation, class of worker etc. We use this task as a stand-in for some task where individuals receive a benefit from being evaluated positively. To get group-calibrated estimates of $\mathbb{P}(r_{i}|\mathcal{D})$, we train a gradient boosting classifier followed by Platt Scaling on the validation subset of the data. We evaluate the EOR criterion and costs on the test subset of these records. Full details for dataset pre-processing and training can be found in Section E.3. Because these rankings are large (up to $\sim 20K$ size), $\pi^{EXP}$ and $\pi^{FS}$ are not computationally tractable. $\pi^{RA}$ performs similarly to $\pi^{PRP}$ and we include it in Figure 13 for completeness.
@@ 7.2.1. How do the ranking policies compare when using learned probability estimates?
To evaluate the two-group EOR algorithm, we first only rank individuals labeled as White and Black or African American. Figure 5 (top) shows that EOR ranking is effective even with estimated probabilities. In particular, while the ranking algorithms only use estimated probabilities, the EOR criterion, and costs are evaluated on the true relevance labels from the test set. Nevertheless, $\pi^{EOR}$ still evaluates $\delta$ close to zero and distributes costs among the stakeholders more evenly than the other baseline policies $\pi^{PRP}$, $\pi^{DP}$, and even $\pi^{\text{unif}},\pi^{TS}$ for a specific ranking with median $\sum_{k=1}^{n}|\delta(\sigma_{k})|$. Additional experiments in Section E.3 further confirm these findings.
@@ 7.3. Amazon Shopping Audit
As the second step of the audit, we use the Markup dataset with logged rankings [^4] https://github.com/the-markup/investigation-amazon-brands and compute $p_{i}$ using the calibrated baseline relevance prediction model. The EOR constraint (6) is averaged over queries for the logged rankings, and the EOR rankings are produced by Algorithm 1. Figure 6 shows that there exists a ranking $\sigma^{EOR}$ that has $\delta(\sigma^{EOR}_{k})$ closer to zero for most prefix $k$. The logged rankings from Amazon’s platform show estimated $\delta(\sigma_{k})$ that are farther away from zero for at least some prefixes of $k$, reflecting a potential favoring of Amazon brand products. A limitation of this analysis is that unlike in a real audit where the auditor has access to the production model of $p_{i}$, our baseline model may be subject to hidden confounding, and thus does not provide conclusive evidence of unfairness. In particular, the production rankings may depend on other features beyond product titles (e.g, product descriptions, bullet points, star ratings, etc.). However, the analysis does demonstrate how the EOR criterion can be used for auditing, if the auditor is given access to the production ranking model to avoid confounding. We provide further details in Section E.4 and our source code with experiment implementation can be found here. [^5] https://github.com/RichRast/Fairness_in_Ranking_under_Disparate_Uncertainty
@@ Acknowledgements.
This research was supported in part by NSF Award IIS-2008139. All content represents the opinion of the authors, which is not necessarily shared or endorsed by their respective employers and/or sponsors. We thank Kate Donahue, Marios Papachristou, Aaron Tucker, Sarah Dean, Luke Wang, Emily Ryu, Ashudeep Singh, Taran Pal Singh, and Woojeong Kim for helpful comments and discussions. We also thank the anonymous reviewers at the Epistemic AI, UAI workshop for helpful feedback.
@@ Appendix B Extended Related Work
Our problem setup involves aggregating candidates from groups and while research on fair rank aggregation appears related, the goal there is much different from our work. In particular, fair rank aggregation achieves maximum consensus accuracy when multiple voters rank all candidates subject to fairness constraints of group exposure (Cachel and Rundensteiner, 2023) or p-fairness (Wei et al., 2022). Work on multi sided fairness (Wang and Joachims, 2021; Burke, 2017) similarly considers diversity constraints or exposure-based formulations. Finally, while (Biega et al., 2018; Singh et al., 2021; Saito and Joachims, 2022; Konstantinov and Lampert, 2021) propose an amortized notion of fairness, our work proposes a non-amortized fairness criterion at every position $k$ of the ranking.
Recently, there has been a growing interest in the study of fairness in rankings under uncertainty. The classical desideratum in this literature studies the relation of group-wise calibration for fairness (Pleiss et al., 2017; Kleinberg et al., 2017; Flores et al., 2016; Chouldechova, 2017; Loi and Heitz, 2022). Our work is orthogonal to this discussion. In particular, we only assume that calibrated probability of relevance is given and instead focus on how differential sharpness of probabilities can cause unfairness. (Singh et al., 2021) introduced an approximate notion of fairness that is violated if the principal ranks candidates that appear more than a certain proportion of their estimated relevance distribution. One way to achieve this in expectation is through randomization of relevances drawn from the predictive posterior distribution. Other works have introduced methods that quantify uncertainty in rankings (Wang and Joachims, 2023) to update and learn better estimates of relevances iteratively (Yang et al., 2023). These works do not consider the unfairness caused due to differential uncertainty between groups and their approach is orthogonal to ours. While methods that reduce uncertainty for all groups are needed, we also need to account for unfairness due to the existing disparate uncertainty that is unfortunately widespread in practical settings.
@@ Appendix C Proofs
Objective
$\text{Maximize}_{X\geq 0}f(X)=\frac{P^{T}X}{nRel(A)+nRel(B)}$
(The objective is equivalent to minimizing the total cost = $1-\frac{P^{T}X}{nRel(A)+nRel(B)}$)
subject to
(10) $X$$\leq$ 11 11
(11) $X^{T}\mathbb{1}$$\leq$ $k$
@@
We define $Q_{A,B}\in^{n}$ where each element of $Q_{A,B}$ is $q_{i}(\mathbb{I}_{A}-\mathbb{I}_{B})_{i}$, $q_{i\in g}=\frac{p_{i}}{nRel(g)}$ and $Q_{A,B}=-Q_{B,A}$.
The first constraint (10) ensures valid values for $X$ (dual variables $\lambda_{i}^{\prime}$ ). The second constraint (11) is for selecting $k$ candidates ( dual variable $\lambda_{k}$ ) and the last two constraints (12) and (13) ensure that the ranking solution is EOR-fair optimal (dual variables $\lambda_{A,B}$, $\lambda_{B,A}$). The Dual LP is formed as follows
Dual
$\text{Minimize}_{\lambda\geq 0}g(\lambda)=\delta(\sigma^{EOR}_{k})(\lambda_{A,B}+\lambda_{B,A})+k\lambda_{k}+\sum_{i=1}^{n}\lambda_{i}^{\prime}$
subject to
(14) $Q_{A,B}(\lambda_{A,B}-\lambda_{B,A})+\lambda_{k}+\lambda^{\prime}\geq\frac{P}{nRel(A)+nRel(B)}$
@@
From (15), (16), (20), (21) , the duality gap between EOR solution and the optimal solution is bounded by
@@
Since $\sum_{i=1}^{k_{A}}q_{i}-\sum_{j=1}^{k_{B}}q_{j}=\frac{nRel(A|\sigma_{k})}{nRel(A)}-\frac{nRel(B|\sigma_{k})}{nRel(B)}$, the lemma follows directly from the definition of $\delta(\sigma_{k})$ in Equation 7 and the EOR ranking principle of choosing the candidate that minimizes $\delta(\sigma_{k})$.
@@
The second equality above is obtained by substituting $\lambda_{k}$ from Equation (17) and the last equality by rearranging. We now consider two cases – for elements not selected and selected by the EOR Algorithm respectively.
@@
In Equation (22), if $\lambda_{B,A}\geq 0$, then $\lambda_{A,B}=0$ and with $p_{i}\leq p_{A}$, $q_{i}\leq q_{A}$ the resultant quantity would be negative, which would result in $\lambda_{i}^{\prime}$ clipped to 00.
@@
In Equation (22), if $\lambda_{A,B}\geq 0$, then $\lambda_{B,A}=0$. We can then substitute $\lambda_{A,B}=\frac{1}{nRel(A)+nRel(B)}\left(\frac{p_{A}-p_{B}}{q_{A}+q_{B}}\right)$ in (22),
@@
In Equation (22), if $\lambda_{B,A}\geq 0$, then $\lambda_{A,B}=0$ and with $p_{i}\geq p_{A}$, $q_{i}\geq q_{A}$ the resultant quantity in (22) would be $\geq 0$, so that $\lambda_{i}^{\prime}\geq 0$.
@@
In Equation (22), if $\lambda_{A,B}\geq 0$, then $\lambda_{B,A}=0$. We can then substitute $\lambda_{A,B}=\frac{1}{nRel(A)+nRel(B)}\left(\frac{p_{A}-p_{B}}{q_{A}+q_{B}}\right)$ in (22),
@@
We now show that $\lambda_{k}\geq 0$. From (17),
@@
If $\lambda_{A,B}\geq 0$ , then $\lambda_{B,A}=0$. Substituting $\lambda_{A,B}=\frac{1}{nRel(A)+nRel(B)}\left(\frac{p_{A}-p_{B}}{q_{A}+q_{B}}\right)$ in (23),
@@
The last inequality follows since each of the terms $p_{A},q_{A},p_{B},q_{B}$ are $\geq 0$. If $\lambda_{B,A}\geq 0$, then $\lambda_{A,B}=0$. By substituting $\lambda_{B,A}=\frac{1}{nRel(A)+nRel(B)}\left(-\frac{p_{A}-p_{B}}{q_{A}+q_{B}}\right)$ in (23), we similarly get $\lambda_{k}\geq 0$.
The two duals $\lambda_{A,B},\lambda_{B,A}$ are $\geq 0$ by their construction in (15), (16). Thus, we have shown that all the constructed dual variables $\lambda\geq 0$. ∎
@@
Using the fact that $\lambda_{i}^{\prime}=0$ for $i>k_{A}$ from Lemma C.2, and substituting $\lambda_{k}$ from Equation (17), we get
@@
Using the fact that $\lambda_{i}^{\prime}\geq 0$ for $i\leq k_{A}$ from Lemma C.2, and substituting $\lambda_{k}$ from Equation (17), $\lambda_{i}^{\prime}$ for $i\leq k_{A}$ in (24), we get
@@
We demonstrate the cost optimality bound proved in Theorem 5.1. Figure 7 shows an example with a ranking produced by Linear Program (LP), Integer Linear Program (ILP), and the EOR algorithm along with the upper bound on the cost computed from the duality gap proved in Theorem 5.1. The example is constructed such that $\mathbb{P}(r_{i}|\mathcal{D})_{i\in A}=[1,0.6,0.5,0.5,0.4,0.1,\cdots 0.1],nRel(A)=4$, and $\mathbb{P}(r_{i}|\mathcal{D})_{i\in B}=[1,0.1\cdots 0.1],nRel(B)=4$. Figure 7 shows that at most prefixes $k$, the EOR cost (in red) is optimal coinciding with the cost from ILP solution (in green) as well as with the LP solution (in blue). Further, when the EOR ranking does not coincide with the LP solution, the upper bound $\frac{2\delta(\sigma^{EOR}_{k})}{nRel(A)+nRel(B)}\left|\frac{p_{A}-p_{B}}{q_{A}+q_{B}}\right|$ is relatively small as is shown by the LP + duality gap (in grey).

### v4

@@ 6. Extension to $G$ Groups
In this section, we discuss the extension of the EOR algorithm beyond two groups. In particular, we consider the general case where a candidate belongs to one of G groups $g\in[1\cdots G]$. From Section 4, we can generalize the cost burden to the principal similar to Eq. (4), taking all the groups into account for the normalization factor as follows
@@
(8) $\delta(\sigma)=\max_{g}\left\{\frac{nRel(g|\sigma)}{nRel(g)}\right\}-\min_{g}\left\{\frac{nRel(g|\sigma)}{nRel(g)}\right\}$
@@
Note that the above selection rule is a strict generalization of Algorithm 1 and it reflects the intuition of minimizing the gap in relevance proportions for all the groups. It can be verified that the runtime complexity with selection rule according to Eqs. (8), (9) for a constant number of groups $G$ is ${O}(n\log{n}+Gn)$. Furthermore, we can extend the cost-approximation guarantee to the multi-group case.
@@
The EOR rankings are cost optimal up to a gap of $\phi\delta(\sigma^{EOR}_{k})$ for $G$ groups, with $\delta(\sigma^{EOR}_{k})$ bounded by $\delta_{max}$, such that,
$\phi$ $=$ $\frac{2}{(G-1)\sum_{g=1}^{G}nRel(g)}\left(\sum_{\{A,B\}}\left|\frac{p_{A}-p_{B}}{q_{A}+q_{B}}\right|\right)~{}~{}~{}\forall k$
@@ 7. Experimental Evaluation
We now evaluate the EOR framework and algorithm empirically and compare against several baselines – namely Demographic (Statistical) Parity ($\pi^{DP}$) (Yang and Stoyanovich, 2016), FAIR Ranking Principle ($\pi^{FS}$) (Zehlike et al., 2017), Probability Ranking Principle ($\pi^{PRP}$) (Robertson, 1977), Thompson Sampling Policy ($\pi^{TS}$) (Singh et al., 2021), Uniform Policy ($\pi^{\text{unif}}$), Disparate Treatment of Exposure ($\pi^{EXP}$) (Singh and Joachims, 2018), and Fair Rank Aggregation ($\pi^{RA}$) (Cachel and Rundensteiner, 2023) with proportional representation of exposure. We discuss implementation details of these baselines in Appendix E.1.
@@ 7.1.1. How does $\pi^{EOR}$ compare against the baselines under varying amounts of disparate uncertainty?
As predicted by theory, $\pi^{EOR}$ maintains low unfairness at all levels of disparate uncertainty, outperforming all the baselines $\pi^{PRP}$, $\pi^{DP}$, $\pi^{TS}$, $\pi^{EXP}$, $\pi^{RA}$, and $\pi^{FS}$. Note that $\pi^{EOR}$ even outperforms the uniform policy $\pi^{\text{unif}}$, since any individual ranking drawn from $\pi^{\text{unif}}$ is likely to be unfair. In terms of effectiveness, the theoretically optimal skyline is given by $\pi^{PRP}$. Across all levels of disparate uncertainty, $\pi^{EOR}$ is at least competitive with the other baselines, indicating that the EOR fairness does not impose a disproportionate cost of fairness for the Principal.
@@ 7.1.2. At which positions in the rankings do the policies incur unfairness?
While the previous table summarized unfairness across the whole ranking, Figure 4 (left) provides more detailed insights into how unfairness accumulates across positions in the ranking. The only method that is systematically fair across all positions $k$ is $\pi^{EOR}$, keeping the unfairness $\delta(\sigma_{k})$ from 4.1 close to zero everywhere in the ranking. The baselines generally start accumulating unfairness towards one group right from the top of the ranking. Their unfairness only decreases once they run out of viable candidates from the group they prefer. The only exception is $\pi^{\text{unif}}$, here for a specific ranking with median $\sum_{k=1}^{n}|\delta(\sigma^{\text{unif}}_{k})|$. However, rankings from $\pi^{\text{unif}}$ tend to stray much further from zero than the $\pi^{EOR}$ ranking. Additional results for the medium and low disparate uncertainty settings in Figure 11 of Appendix E.2 further support these findings.
Figure 4. Left: EOR criterion $\delta(\sigma_{k})$, Middle: group costs according to (3), Right: the principal’s total cost according to (4) of the ranking policies for the synthetic dataset with high disparate uncertainty shown in top right of Table 1. Group A consists of 30 candidates with sharp probabilities with $p_{i}\sim\text{Beta}(1/20,1/20)$. This provides $nRel(A)=14.96$ expected number of relevant candidates. Group B also has similar candidates, in particular, it has 31 candidates, with relatively flat probabilities $p_{i}\sim\text{Beta}(5,5)$, providing $nRel(B)=14.94$ expected number of relevant candidates.
@@ 7.1.3. How do the ranking policies distribute the costs between the stakeholders?
In Figure 4 (middle) we investigate how the ranking policies distribute the cost $c(g|\pi_{k})$ from Eq. (3) between group A and group B. It shows that only $\pi^{EOR}$ has an equal cost to both groups across the whole ranking, which can be seen from the overlapping cost curves for both groups. Furthermore, the cost is substantially lower for both groups than their expected cost under the uniform policy (diagonal line).Figure 4 (right) shows the total cost to the principal, and again $\pi^{EOR}$ is competitive with the baselines.
All other baselines incur substantial disparate costs to the groups, some even worse than the uniform lottery. In particular, $\pi^{DP}$ selects the candidates alternately between the two groups since group sizes are relatively similar, but this results in selecting a higher proportion of relevance from group A because the relevance probabilities are sharper for group A than for B. As a result, the cost burden is higher for group B. $\pi^{TS}$ is fairer than $\pi^{PRP}$, since it randomizes relevant candidates before sorting them in decreasing order of relevance, however being group oblivious, it still places an uneven cost burden.
@@
Figure 5. US Census Dataset: EOR criterion $\delta(\sigma_{k})$ and cost of the ranking policies computed with true relevance labels from the test subset for the US Census dataset. Top: Two groups setting using the White and Black/African American racial groups for the state of Alabama. Bottom: Multiple (four groups) setting using White, Black/African American, Asian, and Other for the state of NY.
@@ 7.2. US Census Survey Data
While the synthetic experiments provide insights into the behavior of ranking policies under varying conditions, we now investigate how far $\pi^{EOR}$ can mitigate unfairness as it arises in real-world datasets where the relevance probabilities $\mathbb{P}(r_{i}|\mathcal{D})$ are learned from data. In particular, we consider the US Census Survey dataset (Ding et al., 2021) for the year 2018 and the state of Alabama and New York, consisting of 22,268 and 103,021 records respectively. The task is to predict whether the income for an individual >\$$>\$50K$ based on features such as educational attainment, occupation, class of worker etc. We use this task as a stand-in for some task where individuals receive a benefit from being evaluated positively. To get group-calibrated estimates of $\mathbb{P}(r_{i}|\mathcal{D})$, we train a gradient boosting classifier followed by Platt Scaling on the validation subset of the data. We evaluate the EOR criterion and costs on the test subset of these records. Full details for dataset pre-processing and training can be found in Section E.3. Because these rankings are large (up to $\sim 20K$ size), $\pi^{EXP}$ and $\pi^{FS}$ are not computationally tractable. $\pi^{RA}$ performs similarly to $\pi^{PRP}$ and we include it in Appendix E.3 for completeness.
@@ 7.2.1. How do the ranking policies compare when using learned probability estimates?
To evaluate the two-group EOR algorithm, we first only rank individuals labeled as White and Black or African American. Figure 5 (top) shows that EOR ranking is effective even with estimated probabilities. In particular, while the ranking algorithms only use estimated probabilities, the EOR criterion, and costs are evaluated on the true relevance labels from the test set. Nevertheless, $\pi^{EOR}$ still evaluates $\delta$ close to zero and distributes costs among the stakeholders more evenly than the other baseline policies $\pi^{PRP}$, $\pi^{DP}$, and even $\pi^{\text{unif}},\pi^{TS}$ for a specific ranking with median $\sum_{k=1}^{n}|\delta(\sigma_{k})|$. Additional experiments in Appendix E.3 further confirm these findings.
@@ 7.3. Amazon Shopping Audit
As the second step of the audit, we use the Markup dataset with logged rankings [^4] https://github.com/the-markup/investigation-amazon-brands and compute $p_{i}$ using the calibrated baseline relevance prediction model. The EOR criterion (7) is averaged over queries for the logged rankings, and the EOR rankings are produced by Algorithm 1. Figure 6 shows that there exists a ranking $\sigma^{EOR}$ that has $\delta(\sigma^{EOR}_{k})$ closer to zero for most prefix $k$. The logged rankings from Amazon’s platform show estimated $\delta(\sigma_{k})$ that are farther away from zero for at least some prefixes of $k$, reflecting a potential favoring of Amazon brand products. A limitation of this analysis is that unlike in a real audit where the auditor has access to the production model of $p_{i}$, our baseline model may be subject to hidden confounding, and thus does not provide conclusive evidence of unfairness. In particular, the production rankings may depend on other features beyond product titles (e.g, product descriptions, bullet points, star ratings, etc.). However, the analysis does demonstrate how the EOR criterion can be used for auditing, if the auditor is given access to the production ranking model to avoid confounding. We provide further details in Figure E.4 and our source code with experiment implementation can be found here. [^5] https://github.com/RichRast/DisparateUncertainty
@@ Acknowledgements.
This research was supported in part by NSF Awards IIS-2008139 and IIS-2312865. All content represents the opinion of the authors, which is not necessarily shared or endorsed by their respective employers and/or sponsors. We thank Kate Donahue, Marios Papachristou, Aaron Tucker, Sarah Dean, Luke Wang, Emily Ryu, Ashudeep Singh, Taran Pal Singh, and Woojeong Kim for helpful comments and discussions. We also thank the anonymous reviewers at the Epistemic AI, UAI workshop for helpful feedback.
@@ Appendix B Extended Related Work
Our problem setup involves aggregating candidates from groups and while research on fair rank aggregation appears related, the goal there is much different. In particular, fair rank aggregation achieves maximum consensus accuracy when multiple voters rank all candidates subject to fairness constraints of group exposure (Cachel and Rundensteiner, 2023) or p-fairness (Wei et al., 2022). Work on multi sided fairness (Wang and Joachims, 2021; Burke, 2017) similarly considers diversity constraints or exposure-based formulations. Finally, while (Biega et al., 2018; Singh et al., 2021; Saito and Joachims, 2022; Konstantinov and Lampert, 2021) propose an amortized notion of fairness, our work proposes a non-amortized fairness criterion at every position $k$ of the ranking.
Recently, there has been a growing interest in the study of fairness in rankings under uncertainty. The classical desideratum in this literature studies the relation of group-wise calibration for fairness (Pleiss et al., 2017; Kleinberg et al., 2017; Flores et al., 2016; Chouldechova, 2017; Loi and Heitz, 2022). Our work is orthogonal to this discussion. In particular, we only assume that calibrated probability of relevance is given and instead focus on how differential sharpness of probabilities can cause unfairness. (Singh et al., 2021) introduced an approximate notion of fairness that is violated if the principal ranks candidates that appear more than a certain proportion of their estimated relevance distribution. One way to achieve this in expectation is through randomization of relevances drawn from the predictive posterior distribution. Other works have introduced methods that quantify uncertainty in rankings (Wang and Joachims, 2023) to update and learn better estimates of relevances iteratively (Yang et al., 2023). These works do not consider the unfairness caused due to differential uncertainty between groups. While methods that reduce uncertainty for all groups are needed, we also need to account for unfairness due to the existing disparate uncertainty that is unfortunately widespread in practical settings.
@@ Appendix C Proofs
(Primal) $\max_{X\geq 0}\quad f(X)$ $=\frac{P^{T}X}{nRel(A)+nRel(B)}$
[…]
(select up to k elements) $X^{T}\mathbb{1}$ $\leq k$
@@
We define $Q_{A,B}\in^{n}$ where each element of $Q_{A,B}$ is $q_{i}(\mathbb{I}_{A}-\mathbb{I}_{B})_{i}$, $q_{i\in g}=\frac{p_{i}}{nRel(g)}$ and $Q_{A,B}=-Q_{B,A}$. Note that the Primal objective is equivalent to minimizing the total cost = $1-\frac{P^{T}X}{nRel(A)+nRel(B)}$.
The first constraint (10) ensures valid values for $X$ (with corresponding dual variables $\lambda_{i}^{\prime}$ ). The second constraint is for selecting $k$ candidates (dual variable $\lambda_{k}$ ) and the last two constraints (11) and (12) ensure that the ranking solution is EOR-fair optimal (dual variables $\lambda_{A,B}$, $\lambda_{B,A}$). The Dual LP is formed as follows
(Dual) $\min_{\lambda\geq 0}$ $\quad g(\lambda)=\delta(\sigma^{EOR}_{k})(\lambda_{A,B}+\lambda_{B,A})+k\lambda_{k}+\sum_{i=1}^{n}\lambda_{i}^{\prime}$
(13) s.t. $\quad Q_{A,B}(\lambda_{A,B}-\lambda_{B,A})+\lambda_{k}+\lambda^{\prime}\geq\frac{P}{nRel(A)+nRel(B)}$
@@
From Eqs. (14), (15), (19), (20) , the duality gap between EOR solution and the optimal solution is bounded by
@@
Since $\sum_{i=1}^{k_{A}}q_{i}-\sum_{j=1}^{k_{B}}q_{j}=\frac{nRel(A|\sigma_{k})}{nRel(A)}-\frac{nRel(B|\sigma_{k})}{nRel(B)}$, the lemma follows directly from the definition of $\delta(\sigma_{k})$ in Eq. (7) and the EOR ranking principle of choosing the candidate that minimizes $\delta(\sigma_{k})$.
@@
The second equality above is obtained by substituting $\lambda_{k}$ from Eq. (16) and the last equality by rearranging. We now consider two cases – for elements not selected and selected by the EOR Algorithm respectively.
@@
In Eq. (21), if $\lambda_{B,A}\geq 0$, then $\lambda_{A,B}=0$ and with $p_{i}\leq p_{A}$, $q_{i}\leq q_{A}$ the resultant quantity would be negative, which would result in $\lambda_{i}^{\prime}$ clipped to 00.
@@
In Eq. (21), if $\lambda_{A,B}\geq 0$, then $\lambda_{B,A}=0$. We can then substitute $\lambda_{A,B}=\frac{1}{nRel(A)+nRel(B)}\left(\frac{p_{A}-p_{B}}{q_{A}+q_{B}}\right)$ in Eq. (21),
@@
In Eq. (21), if $\lambda_{B,A}\geq 0$, then $\lambda_{A,B}=0$ and with $p_{i}\geq p_{A}$, $q_{i}\geq q_{A}$ the resultant quantity in (21) would be $\geq 0$, so that $\lambda_{i}^{\prime}\geq 0$.
@@
In Eq. (21), if $\lambda_{A,B}\geq 0$, then $\lambda_{B,A}=0$. We can then substitute $\lambda_{A,B}=\frac{1}{nRel(A)+nRel(B)}\left(\frac{p_{A}-p_{B}}{q_{A}+q_{B}}\right)$ in (21),
@@
We now show that $\lambda_{k}\geq 0$. From Eq. (16),
@@
If $\lambda_{A,B}\geq 0$ , then $\lambda_{B,A}=0$. Substituting $\lambda_{A,B}=\frac{1}{nRel(A)+nRel(B)}\left(\frac{p_{A}-p_{B}}{q_{A}+q_{B}}\right)$ in Eq. (22),
@@
The last inequality follows since each of the terms $p_{A},q_{A},p_{B},q_{B}$ are $\geq 0$. If $\lambda_{B,A}\geq 0$, then $\lambda_{A,B}=0$. By substituting $\lambda_{B,A}=\frac{1}{nRel(A)+nRel(B)}\left(-\frac{p_{A}-p_{B}}{q_{A}+q_{B}}\right)$ in Eq. (22), we similarly get $\lambda_{k}\geq 0$.
The two duals $\lambda_{A,B},\lambda_{B,A}$ are $\geq 0$ by their construction in Eqs. (14), (15). Thus, we have shown that all the constructed dual variables $\lambda\geq 0$. ∎
@@
Using the fact that $\lambda_{i}^{\prime}=0$ for $i>k_{A}$ from Lemma C.2, and substituting $\lambda_{k}$ from Eq. (16), we get
@@
Using the fact that $\lambda_{i}^{\prime}\geq 0$ for $i\leq k_{A}$ from Lemma C.2, and substituting $\lambda_{k}$ from (16), $\lambda_{i}^{\prime}$ for $i\leq k_{A}$ in (23), we get
@@
We demonstrate the cost optimality bound proved in Theorem 5.1 in Figure 7 that shows an example with a ranking produced by Linear Program (LP), Integer Linear Program (ILP), and the EOR algorithm along with the upper bound on the cost computed from the duality gap proved in Theorem 5.1. The example is constructed such that $\mathbb{P}(r_{i}|\mathcal{D})_{i\in A}=[1,0.6,0.5,0.5,0.4,0.1,\cdots 0.1],nRel(A)=4$, and $\mathbb{P}(r_{i}|\mathcal{D})_{i\in B}=[1,0.1\cdots 0.1],nRel(B)=4$. Figure 7 shows that at most prefixes $k$, the EOR cost (in red) is optimal coinciding with the cost from ILP solution (in green) as well as with the LP solution (in blue). Further, when the EOR ranking does not coincide with the LP solution, the upper bound $\frac{2\delta(\sigma^{EOR}_{k})}{nRel(A)+nRel(B)}\left|\frac{p_{A}-p_{B}}{q_{A}+q_{B}}\right|$ is relatively small as is shown by the LP + duality gap (in grey).